OpenStudy (anonymous):

The Cosine theorem: c^2=a^2-2*a*b*cos(C). How to find the angles if you know all sides?

OpenStudy (anonymous):

unless it's a specific type of triangle, you can't...i think...

OpenStudy (anonymous):

i stand corrected...it seems you can...

OpenStudy (anonymous):

I found the angle C but i can't find the others: C=acos((c^2-a^2-b^2)/-2*a*b)

OpenStudy (anonymous):

from that you can use sine law \[\sin A = \frac{a\sin C}{c}\]

OpenStudy (anonymous):

since sine law states \[\frac{\sin A}a = \frac{\sin C}c\] or are you not allowed to use that?

OpenStudy (anonymous):

Yhea, Sin law is good. But does it work with the Cosine theorem?

OpenStudy (anonymous):

if memory serves right... the three equations of cosine law constains 2 sides and 1 angle...so i think it's possible

OpenStudy (anonymous):

C=acos((c^2-a^2-b^2)/(-2*a*b) B=acos((b^2-a^2-c^2)/(-2*a*c) A=acos((a^2-b^2-c^2)/(-2*b*c) Ok, i found the solution.

OpenStudy (anonymous):

congratulations

OpenStudy (anonymous):

oh, i got 2 medals! thanks.

OpenStudy (anonymous):

users who solve their own questions tend to get medals